OpenStudy (anonymous):

8/9-x/3=2x-2... solve

OpenStudy (vishweshshrimali5):

i can help

OpenStudy (vishweshshrimali5):

first of all multiply by 9 both sides

OpenStudy (anonymous):

ok

OpenStudy (vishweshshrimali5):

you will get

OpenStudy (anonymous):

will you write out the steps for ,me please

OpenStudy (vishweshshrimali5):

8/9-x/3=2x-2 => 8 - 3x = 18x - 18 => 18x + 3x = 8 + 18 => 21 x = 26 => x = 26/21 This is your required answer

OpenStudy (anonymous):

thank yoi

OpenStudy (vishweshshrimali5):

Hey mention not friend just give me a medal if you are satisfied